2. A boy with a mass of 60kg runs at a speed of 3m / s behind a cart weighing 160kg, which moves in the same direction at a speed of 1m / s.
When they reach the cart, it is thrown into it. Find the size and direction of the cart as the boy jumps on it. Friction is not taken into account.
1
Expert's answer
2021-02-12T06:10:19-0500

From the law of conservation of momentum, we get:


"m_1v_1+m_2v_2=(m_1+m_2)v,""v=\dfrac{m_1v_1+m_2v_2}{m_1+m_2},""v=\dfrac{60\ kg\cdot3\ \dfrac{m}{s}+160\ kg\cdot1\ \dfrac{m}{s}}{60\ kg+160\ kg}=1.54\ \dfrac{m}{s}."

The sign plus means that the cart and boy moves in the same direction (to the right).